OpenStudy (anonymous):

solve: log base 2 (3t +1) = 4

OpenStudy (amistre64):

log2(3t +1) = 4 2^(log2(3t+1)) = 2^4 3t+1 = 2^4 3t = 16 - 1 t = 15/3 t=5

OpenStudy (anonymous):

gracias :)

OpenStudy (amistre64):

de nada

OpenStudy (anonymous):

are you still on?